A health specialist wants to determine the average number of hours a person

exercises in a day during the quarantine period. She found out that the mean number

of hours a person exercises in a day during the quarantine period is 80 minutes. A

random sample of 29 persons were surveyed and found that their mean is 65 minutes

and a standard deviation of 10 minutes. Test the hypothesis at 2% level of significance

and assume that the population is normally distributed.


1
Expert's answer
2022-06-12T15:45:53-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=80"

"H_1:\mu\not=80"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.02," "df=n-1=28" and the critical value for two-tailed test is "t_c =2.46714."

The rejection region for this two-tailed test is "R = \{t:|t|>2.46714\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{65-80}{10/\sqrt{29}}=-8.0777"


Since it is observed that "|t|=8.0777>2.46714=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=28" degrees of freedom, "t=-8.0777" is "p=0," and since "p=0<0.02=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 80, at the "\alpha = 0.02" significance level.